A manufacturer knows that their items have a normally distributed lifespan, with a mean of 7.5 years, and standard deviation of 1.8 years.

If you randomly purchase one item, what is the probability it will last longer than 12 years?
Answer #1

Solution :

Given that,

mean = = 7.5

standard deviation = = 1.8   

P(X>12 ) = 1 - P[(X- ) / < (12 - 7.5) /1.8 ]

= 1 - P(z < 2.5)

Using z table

= 1 - 0.9938   

=0.0062